The area of a square is (4x^2-12x+9) sq.cm.The width of a rectangle is five cm greater than a side of the square while the length of the rectangle is twice the side of the square.What is the area, in sq.cm. of the rectangle? please show me the steps.
A.16x^2-12x-14
B.8x^2+10x-12
C.4x^2-8x-16
D.4x^2+2x-2
E.None of these

Answer by flame8855(424)

4x^2-12x+9 is the area of the square
4x^2-12x+9= s^2 = (2x-3)^2
so the side of the square is 2x-3
the width = 2x-3+5=2x+2
the length = 2(2x-3)=4x-6
area= L*W= (2x+2)(4x-6)= 8x^2-12x+8x-12= 8x^2-4x-12
so none of the above